1.Calculate the rfm of gasA given that the time taken for equal volumes of oxygen and gas A to diffuse through a hole is 20sec and 24 sec respectively
2.A given volume of ozone O3 ,diffused from a certain apparatus in 96 sec.calculte the time taken by an equal volume of carbon (iv)oxide to diffuse under the same condition
3.given that the time taken for equal volume of gas A and gas B to diffuse through the same hole is 20 and 24 , calculate the molecular mass of gas B if the molecular mass of A is 32
1
Expert's answer
2020-04-19T07:10:23-0400

1. R1/R2=√M2/M1

M2=(R1/R2)^2 x M1= (24/20)^2 x 32= 46.08.

2. R1/R2=√M2/M1

V2=(R1/R2)^2 x V1= (96/96)^2 x V1= V1.

3. R1/R2=√M2/M1

M2=(R1/R2)^2 x M1= (24/20)^2 x 32= 46.08.